I am 12 years old my sister is 5 years old. In how many years will I be twice as old as my sister
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52803)

.

Let x = in how many years . . . 


Then you have this equation to determine x


    12 + x = 2*(5+x)


Simplify and solve for x


    12 + x = 10 + 2x

    12 - 10 = 2x - x

     2      = x.


ANSWER.  In 2 years.


Answer by greenestamps(13200)



The difference between your ages is 7 years. For the sake of a problem in mathematics, that difference will always be 7 years.

So you will be twice as old as your sister when she is 7 and you are 14.

Since you are now 12, that will be 2 years from now.